OpenStudy (anonymous):

Why does this simplify to what it is: (as seen below) [I know the basic fractions to power prin.]

OpenStudy (anonymous):

x^(0.5)^(0.5)^(0.5) = x^2 ^ (-1/sqrt(2))

zepdrix (zepdrix):

Are we talking about this?\[\Large\bf\sf x^{1/2^{1/2^{1/2}}}\]Or this?\[\Large\bf\sf \left(\left(x^{1/2}\right)^{1/2}\right)^{1/2}\]

OpenStudy (anonymous):

Don't know, the above is how it was inter-pretended.

OpenStudy (anonymous):